### v2

# FoMEMO: Towards Foundation Models for Expensive Multi-objective Optimization

Abstract
Expensive multi-objective optimization is a prevalent and crucial concern in many real-world scenarios, where sample-efficiency is vital due to the limited evaluations to recover the true Pareto front for decision making. Existing works either involve rebuilding Gaussian process surrogates from scratch for each objective in each new problem encountered, or rely on extensive past domain experiments for pre-training deep learning models, making them hard to generalize and impractical to cope with various emerging applications in the real world. To address this issue, we propose a new paradigm named FoMEMO (Foundation Models for Expensive Multi-objective Optimization), which enables the establishment of a foundation model conditioned on any domain trajectory and user preference, and facilitates fast in-context optimization based on the predicted preference-wise aggregated posteriors. Rather than accessing extensive real-world domain experiments for training, we demonstrate that pre-training the foundation model with a diverse set of hundreds of millions of synthetic data can lead to superior generalization and optimization performance to unknown problems, without necessitating any subsequent model training or updates in the following optimization process.
Machine Learning, ICML
1 Introduction
Optimizing multiple competitive objectives simultaneously is ubiquitous in many real-world applications, such as neural architecture search (Ying et al., 2019), engineering design (Tanabe and Ishibuchi, 2020) and scientific discovery (Dara et al., 2022). The primary difficulty in these scenarios stems from the black-box nature of real-world evaluations, which are often costly and time-consuming. When decision-makers have access to only a limited number of evaluations, it becomes challenging to accurately recover the unknown true Pareto fronts.
To achieve sample-efficiency, traditional multi-objective Bayesian optimization (MOBO) typically employs Gaussian Process (GP) surrogates (Williams and Rasmussen, 1995) to approximate objective functions, and then optimize acquisition functions to generate candidate solutions to update the surrogates iteratively in an online manner (Daulton et al., 2020, 2021). However, these methods typically need to rebuild surrogates from scratch for each objective when encountering a new problem, and the efficiency is greatly limited by the cost of GP training and inference (Williams and Rasmussen, 2006). To enhance the generalization and scalability in downstream optimization, some research focuses on pre-training deep learning models in offline scenarios (Xue et al., 2024; Yuan et al., 2024). These approaches typically assume accessing to a large amount of real-world experimental data is available for training. However, acquiring large datasets in practical applications, especially in unknown or emerging fields, is often challenging or even impossible. While recent work has explored pre-training with synthetic functions in the online setting (Hung et al., 2025), its architecture still depends heavily on the repetitive GP modeling. Furthermore, the reliance on computationally intensive hypervolume-based rewards (Zitzler and Thiele, 2002) imposes a significant bottleneck on training scalability, thereby restricting the diversity of the training data and hindering its generalization to unseen scenarios.
In real-world contexts, a wide range of multi-objective optimization problems arises daily across various fields, each exhibiting unique characteristics and demand rapid solutions. Given the limitations of existing methodologies, a compelling research question arises:
Can we devise a universal method that can efficiently address a wide range of emerging real-world multi-objective optimization problems?
In this paper, we attempt to answer this question by exploring a new paradigm named FoMEMO. It enables us to establish a foundation model conditioned on an arbitrary domain trajectory and user preference, and facilitates fast in-context optimization based on the predicted preference-wise aggregated posteriors. What distinguishes our approach from existing methods lies in two aspects. Firstly, unlike traditional MOBO methods that require rebuilding a surrogate from scratch for each objective in newly encountered problems and necessitate frequent, costly model updates during optimization, our approach involves pre-training a foundation model only once, which can be directly adapted to arbitrary scenarios throughout the entire optimization process in an in-context manner. This means that by simply providing domain trajectories and user preferences as contextual information, the pre-trained model can achieve few-shot generalization in unknown scenarios to initiate fast optimization without any subsequent model training or updates. Secondly, rather than accessing expensive real-world data from extensive past experiments for training, we continuously sample a large amount of diverse synthetic data from a computationally efficient sampling mechanism during pre-training. By simulating a large set of potential scenarios that may exist in the real world, the foundation model emerges with the ability to learn to predict the aggregated posteriors conditioned on any contexts.
To the best of our knowledge, this is the first attempt to establish a foundation model for addressing expensive multi-objective optimization problems in an in-context manner. The main contributions of this paper are as follows:
  • We pre-train a foundation model using hundreds of millions of sampled synthetic data, enabling it to learn to predict aggregated posteriors conditioned on arbitrary domain trajectories and user preferences. This approach eliminates the need for access to extensive expensive experimental data from the real world during training, allowing for efficient and adaptable optimization across diverse scenarios.
  • Based on the predicted aggregated posteriors, we develop preference-based and preference-free acquisition functions that can be quickly optimized to generate candidate solutions in an in-context manner, without requiring any subsequent model updates.
  • We evaluate the proposed method on a diverse set of synthetic problems and real-world applications. Experimental results demonstrate that the method exhibits remarkable adaptability and generalization capabilities across a wide range of unseen problems. In comparison to existing methods, our approach consistently achieves superior optimization performance and high efficiency in the majority of cases.
2 Related Work
2.1 Multi-objective Bayesian Optimization
Multi-objective Bayesian optimization (MOBO) typically employs GP models and acquisition functions to identify Pareto optimal solutions within a constrained evaluation budget. A popular category of MOBO is based on decomposition, wherein the original MOP is converted into several single-objective sub-problems that are easier to solve (e.g., ParEGO (Knowles, 2006), TS-TCH (Paria et al., 2020) and MOEA/D-EGO (Zhang et al., 2009)). Besides decomposition, indicator-based approaches provide an integrated performance measure across all objectives under uncertainty. Prominent examples include EHVI (Emmerich et al., 2006), which generalizes the Expected Improvement (EI) function (Jones et al., 1998) to multi-objective settings, along with its modern variants: the parallelizable qEHVI (Wada and Hino, 2019; Daulton et al., 2020) and the noise-robust qNEHVI (Daulton et al., 2021). Another representative class stems from information theory, focusing on selecting points that maximize information gain regarding the Pareto set or front (e.g., PESMO (Hernández-Lobato et al., 2016), MESMO (Belakaria et al., 2019) and JES (Tu et al., 2022; Hvarfner et al., 2022). In addition to model management and acquisition functions design, various batch sample selection strategies have also been developed in parallel setting (e.g., TSEMO (Bradford et al., 2018), USEMO-EI (Belakaria et al., 2020) and DGEMO (Konakovic Lukovic et al., 2020)). While effective, these traditional MOBO methods typically require training individual models for each objective of every new problem from scratch. Furthermore, the substantial computational overhead associated with GP training and inference significantly hinders their scalability.
2.2 Learning-based Expensive Optimization
To enhance the generalization and scalability in downstream optimization, several approaches in the offline context involves training a deep neural network (DNN) on a static dataset to serve as an oracle (Trabucco et al., 2021; Yu et al., 2021; Chen et al., 2023). This is followed by applying search methods, such as gradient descent or evolutionary algorithms, to generate a set of candidate solutions that are potentially superior to those present in the dataset (Xue et al., 2024; Yuan et al., 2024). However, these approaches typically require a large amount of training data from expensive experiments, which is often impractical for various emerging applications in the real-world. Moreover, they often necessitate the training of dedicated models for specific problems, which limits their ability to generalize to unseen scenarios. To circumvent the need for extensive real-world training data, inspired by meta-learning frameworks in single-objective BO (Volpp et al., 2019; Hsieh et al., 2021; Maraval et al., 2024), BOFormer (Hung et al., 2025) has recently explored pre-training policies via reinforcement learning on synthetic data in multi-objective context. However, to predict acquisition function values, its architecture remains heavily dependent on iterative GP modeling to generate posterior distributions. Furthermore, the computationally intensive nature of hypervolume-oriented rewards imposes a significant bottleneck on training scalability, thereby restricting the diversity of the pre-training corpus and hindering its generalization to unseen complex problems.
2.3 Universal Regression Based on Transformers
Recently, Transformers (Vaswani et al., 2017) have become the preferred architecture for deep learning and foundation models. As model sizes and training data scaling, the models exhibit remarkable in-context learning (ICL) capabilities (Brown et al., 2020): they can make accurate predictions quickly in new tasks when prompted with only a few training examples, without any parameter update. Several research has demonstrated that Transformers can learn in-context simple logistic regression algorithms (Garg et al., 2022; Akyürek et al., 2022). Beyond predicting labels, Transformer Neural Processes (Nguyen and Grover, 2022) enhance the prediction with uncertainty as an alternative framework for uncertainty-aware meta learning. Prior-data Fitted Networks (PFNs) (Müller et al., 2021) further show that Transformers can even learn to perform Bayesian inference with ICL. The developed universal regression models not only provide new insights into various model-based prediction topics (Hollmann et al., 2022, 2025; Dooley et al., 2023; Song et al., 2024a), but also promote the regress-then-optimize framework towards a generalized optimization paradigm in both numerical (Müller et al., 2023; Nguyen et al., 2023; Rakotoarison et al., 2024) and even linguistic space (Nguyen et al., 2024; Tan et al., 2025). Building on this potential, we aim to develop a general optimization method for multi-objective scenarios, aligning more closely with real-world challenges that involve multi-criteria outputs.
3 Preliminaries
3.1 Expensive Multi-objective Optimization
In this paper, we address the following multi-objective optimization problem (MOP):
$\min_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{f}(\boldsymbol{x})=(f_{1}(\boldsymbol{x}),f_{2}(\boldsymbol{x}),\cdots,f_{m}(\boldsymbol{x})),$ (1)
where $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{d})^{T}$ is the feature vector in the decision space $\mathcal{X}\subset\mathbb{R}^{d}$, and $\boldsymbol{f}:\mathcal{X}\rightarrow\mathbb{R}^{m}$ is the objective vector containing $m$ continuous objective functions. We consider the scenario in which all the objective functions are expensive to evaluate, with no known analytical expressions or gradient information. Typically, there is no single solution that can simultaneously achieve the optimal for all objectives. Instead, the goal is to find a set of solutions with Pareto optimal objective trade-offs. We say an objective vector $\boldsymbol{f}(\boldsymbol{x})$ Pareto dominates another $\boldsymbol{f}(\boldsymbol{x}^{\prime})$, denoted as $\boldsymbol{f}(\boldsymbol{x})\prec\boldsymbol{f}(\boldsymbol{x}^{\prime})$ if $f_{i}(\boldsymbol{x})\leq f_{i}(\boldsymbol{x}^{\prime}),\forall i\in\{1,\ldots,m\}$ and there exists $j\in\{1,\ldots,m\}$ such that $f_{j}(\boldsymbol{x})<f_{j}(\boldsymbol{x}^{\prime})$. A solution $\boldsymbol{x}^{*}\in\mathcal{X}$ is Pareto optimal if there is no $\boldsymbol{x}$ such that $\boldsymbol{f}(\boldsymbol{x})\prec\boldsymbol{f}(\boldsymbol{x}^{*})$. The set of all Pareto optimal solutions is called the Pareto set (PS), and the image of PS in the objective space is called the Pareto front (PF).
3.2 Scalarization for Multiple Objectives
Scalarization is a straightforward and effective technique for transforming multiple objectives into a single metric. A classic approach is based on decomposition, in which multiple objectives are aggregated using a family of scalarization functions $s_{\boldsymbol{\lambda}}(\boldsymbol{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ parameterized by a set of preferences $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$. We consider the popular Tchebycheff scalarization function:
$s_{\boldsymbol{\lambda}}^{\text{tch}}(\boldsymbol{x})=\max_{1\leq i\leq m}\{\lambda_{i}(y_{i}(\boldsymbol{x})-z_{i}^{*})\},$ (2)
where $y_{i}(\boldsymbol{x})$ is the $i$-th observation value of the corresponding true function $f_{i}(\boldsymbol{x})$ at the input $\boldsymbol{x}$, $\boldsymbol{z}^{*}=(z_{1}^{*},\cdots,z_{m}^{*})$ is the ideal point (lower bound) in the objective space. A promising property behind Tchebycheff scalarization function is that we can identify each Pareto optimal solution by optimizing the scalarization function with a specific but unknown preference (Choo and Atkins, 1983):
$\boldsymbol{x}^{*}=\arg\min_{\boldsymbol{x}\in\mathcal{X}}s_{\boldsymbol{\lambda}}^{\text{tch}}(\boldsymbol{x}).$ (3)
In our work, we use the negative of the scalarization function as the aggregation target $g_{\boldsymbol{\lambda}}(\boldsymbol{x})=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})$ for the convenience of model training and inference. Thus, we wish to maximize $g_{\boldsymbol{\lambda}}(\boldsymbol{x})$ given each preference vector $\boldsymbol{\lambda}$.
Another representative scalarization method is the hypervolume (HV) indicator (Zitzler and Thiele, 2002), which is monotonic to the Pareto dominance relation. The hypervolume of a finite approximate Pareto front $\mathcal{P}$ is defined as:
$\text{HV}(\mathcal{P})=\boldsymbol{\Lambda}_{m}(\{\boldsymbol{v}\in\mathbb{R}^{m}|\exists\boldsymbol{p}\in\mathcal{P}:\boldsymbol{p}\prec\boldsymbol{v}\prec\boldsymbol{r}\}),$ (4)
where $\boldsymbol{\Lambda}_{m}$ denotes the m-dimensional Lebesgue measure, and $\boldsymbol{r}\in\mathbb{R}^{m}$ is a pre-defined reference point.
3.3 Bayesian Optimization
Bayesian optimization (BO) (Shahriari et al., 2015; Frazier, 2018; Garnett, 2023) is a sample-efficient technique for solving expensive black-box optimization problems. BO utilizes probabilistic surrogate models, typically Gaussian processes (GPs) (Williams and Rasmussen, 2006), to provide an prediction with calibrated uncertainty for the latent functions $\boldsymbol{f}$ based on the true observed data $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$. To suggest the candidate solutions for next evaluations, BO defines and optimizes an acquisition function $\alpha:\mathcal{X}\mapsto\mathbb{R}$, which quantifies the utility value of evaluating a new set of solutions based on the predictive distribution of the surrogate models, effectively balancing exploration and exploitation.
4 Methodology
Figure 1: Overview of FoMEMO framework, including synthetic pre-training stage (left) and in-context optimization stage (right). During synthetic pre-training, the foundation model is fed with a large set of synthetic data as the context inputs, which include the trajectory pairs $D_{n}=\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, the query input $\boldsymbol{x}$ with its aggregation target $g=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})$ masked, and the corresponding artificial preference $\boldsymbol{\lambda}$ sampled on the simplex. The model parameterized by $\boldsymbol{\theta}$ is trained once only to predict the aggregated posterior distributions $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$ as the output conditioned on the contextual information. During the in-context optimization stage, users can readily address new problems by simply providing evaluated trajectories from arbitrary unseen real-world applications, along with any potential user preferences (if applicable) as prompt inputs for the foundation model. Using the predicted aggregated posteriors as the basis, various acquisition functions can be naturally derived to enable fast in-context optimization. These acquisition functions can be efficiently optimized to suggest the next candidate within the utility of interest, free from any additional model training or updates throughout the entire optimization process.
4.1 Framework
Instead of building a separate surrogate model from scratch for each objective function in every new problem encountered, the core concept of our method is to develop a universal foundation model. Our foundation model is pre-trained once only using a vast corpus of synthetic samples and preferences as context information, and outputs a large set of preference-wise aggregated posterior distributions for fast in-context optimization. We present the details of our framework in Figure 1, which includes a synthetic pre-training stage on the left and an in-context optimization stage on the right.
Specifically, in the synthetic pre-training stage, rather than accessing expensive real-world experiments, we implement a data sampling mechanism that allows us to heavily sample from it to generate a large set of synthetic data as the context inputs, including the trajectory pairs $D_{n}=\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, the query input $\boldsymbol{x}$ with the corresponding aggregation target $g=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})$ masked, and the artificial preference $\boldsymbol{\lambda}$ sampled on the simplex, constructing a supervised learning problem for training the foundation model efficiently. We repeatedly synthesize training data as above and optimize the model’s parameters $\boldsymbol{\theta}$ to predict the aggregated posterior distributions $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$ conditioned on the contextual inputs. We will present the details of data generation in Section 4.2.
Once the pre-training is completed, during the in-context optimization stage, users can tackle new problems by simply providing the evaluated trajectory from arbitrary unseen real-world applications, as well as the user’s potential preferences (if any) as prompt inputs for the foundation model. Based on the aggregated posteriors predicted by the foundation model in a single neural network forward pass, various acquisition functions can be easily derived to initiate fast in-context optimization to generate the next candidate for expensive evaluation for the utility of interest, without the need for any additional model training or updates during the optimization process.
4.2 Data Generation
While the pre-training distribution should ideally align with that of real-world functions, capturing such a distribution is often challenging due to limited empirical data. To bridge this gap, we implement a data generation mechanism designed to simulate a vast array of surrogate scenarios that the model may encounter. This approach enables the creation of a diverse yet computationally efficient dataset, bolstering the model’s generalization capacity and enabling rapid adaptation to complex, unseen real-world optimization tasks.
Our data generation mechanism consists of three steps. Firstly, we sample general parameters for each set of the data, including the number of features $d$, the number of objectives $m$ and the length of trajectory $n$. Secondly, in order to simulate the multitude of implicit mappings that may exist in the real world, we sample a variety of synthetic functions from Gaussian processes (GPs). The choice of GPs is motivated by their cost-effectiveness and flexibility. Sampling from a GP prior is an easy and cost-effective method that enables the generation of a substantial number of latent mappings at a low cost. Besides, we can easily control the diversity of the generated functions by adjusting the kernel parameters. We employ a GP with the RBF kernel as it is a universal approximator (Micchelli et al., 2006) to encompass a wide range of unknown functions, which means in theory if trained long enough the foundation model will encounter synthetic functions that closely resemble true objective functions, enabling effective generalization. In the third step, we sample the observed trajectory $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$ together with an arbitrary query pair $(\boldsymbol{x},\boldsymbol{y})$ with $d$-dimensional features and $m$-dimensional objectives from each sampled function. Since the aggregation target $g$ is determined only by $\boldsymbol{x}$ given $\boldsymbol{\lambda}$, after sampling artificial preferences from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$, we can calculate the aggregation targets efficiently for each query input given arbitrary preferences. We synthesize hundreds of millions of training data using the aforementioned method, and mask the query aggregation targets to formulate a supervised prediction problem. This processing allows the foundation model to learn to effectively predict the preference-wise posterior distribution $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$.
4.3 Model Architecture
Among the model architectures capable of learning uncertainty-calibrated predictions directly from data, we utilize the Prior-Data Fitted Networks (PFNs) (Müller et al., 2021) as the base framework, as it provides a flexible paradigm for approximating the posterior distribution with arbitrary shape in a principled manner (Nagler, 2023). PFNs are based on a Transformer encoder that does not incorporate positional encodings, allowing the model to process data regardless of the order in which the input elements are presented. We use a variable-dimensional encoder to encode each input $\boldsymbol{x}$ and observation $\boldsymbol{y}$ into a token, allowing the model to adapt to cross-dimensional feature and objective spaces. An attention mask is implemented to enable each trajectory point can only attend to each other and the query point can only attend to the trajectory points. We also encode the sampled preference to inform the model to learn to predict the posterior distribution conditioned on which preference. To approximate the complex posterior distribution in high-dimensional feature and objective spaces, PFNs employ a piecewise constant distribution in the regression head. This approach converts the continuous distribution into a discrete distribution by utilizing multiple intervals, thereby transforming the regression task into a classification task to enhance the robustness during training. The detailed implementation can be referred to (Müller et al., 2021). To account for the distributional shifts in aggregated posteriors across varying objective dimensions, we propose an objective-aware regression head. This architecture dynamically routes the input to a specialized regressor tailored to the specific objective dimension of the problem at hand. By leveraging this dimension-specific mapping, the model can more accurately capture the differences of the aggregated posteriors with higher fidelity, thereby significantly enhancing its predictive precision in unseen landscapes across varying objective dimensions. Please refer to Appendices A.1 and C for detailed model configurations and the ablation analyses with different settings.
4.4 Training Details
For any set of synthetic data generated from the sampling mechanism $p(D)$ presented in Section 4.2, we optimize the following loss function, which is the cross-entropy between the held-out aggregation target and the model prediction:
$\mathbb{E}_{(\boldsymbol{x},y)\cup D_{n}\sim p(D),\boldsymbol{\lambda}\sim\Delta,g=-s_{\boldsymbol{\lambda}}(\boldsymbol{x})}[-\log q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})].$ (5)
We continuously synthesize data through extensive sampling and optimize the model parameters to minimize the loss function, thereby progressively approaching the true posterior distribution $p(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$.
Our final foundation model uses the PFNs architecture with 12 layers of Transformer, an embedding size of 512, a hidden size of 1024 in feed-forward layers, and 4-head attention, resulting in a total of 26.81M parameters. During training, we randomly sample each dataset from a large data space with feature dimensions ranging from 1 to 30, objective dimensions from 1 to 6, and trajectory lengths from 1 to 256. We sample a large variety of functions from GP. Specifically, we randomly sample the length scale kernel parameter $l\sim\Gamma(\alpha=3.0,\beta=6.0)$ for each feature and add a random observation noise $\epsilon\sim\mathcal{N}(0,10^{-4})$ to ensure function diversity. We do not sample the output scale parameter but instead set it to 1.0, as the predicted aggregation target will be normalized during inference. The final model is trained using 500 epochs and 2048 steps in each epoch with a batch size of 128 datasets, resulting more than 130 million datasets in total. The loss function is optimized using Adam optimizer (Kingma and Ba, 2014) with linear warmup and cosine annealing (Loshchilov and Hutter, 2016) using a learning rate of $5\times 10^{-5}$. The entire training process is conducted on 8 NVIDIA L20 GPUs and takes approximately 35 hours. We perform the pre-training only once and use the same model in all experiments in this paper.
4.5 Generating Candidates
Benefiting from the preference-wise aggregated posterior as the output basis, modeling the distribution of infinite sub-problems becomes possible. Acquisition functions such as random scalarization and multi-objective indicator based on scalarization can be well adapted directly. We have developed two types of acquisition functions including preference-based and preference-free in our work.
Preference-based
We firstly derive the preference-based expected improvement (EI) and upper confidence bound (UCB) acquisition functions based on the single-objective scenario (Müller et al., 2023):
$\alpha_{\text{EI}}(\boldsymbol{x};\boldsymbol{\lambda})=\mathbb{E}_{q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})}[[g_{\boldsymbol{\lambda}}(\boldsymbol{x})-g_{\boldsymbol{\lambda}}^{*}]_{+}],$ (6)
$\alpha_{\text{UCB}}(\boldsymbol{x};\boldsymbol{\lambda})={\mu}_{\boldsymbol{\lambda}}(\boldsymbol{x})+\beta{\sigma}_{\boldsymbol{\lambda}}(\boldsymbol{x}),$ (7)
where $g_{\boldsymbol{\lambda}}^{*}$ is the best value of the aggregation $g_{\boldsymbol{\lambda}}(\boldsymbol{x})$ given an arbitrary preference $\boldsymbol{\lambda}$. ${\mu}_{\boldsymbol{\lambda}}(\boldsymbol{x})$ and ${\sigma}_{\boldsymbol{\lambda}}(\boldsymbol{x})$ are the mean and standard deviation of the aggregated posterior $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$, respectively. $\beta$ is a constant, which we set to 1.0 in this paper. As inference the acquisition functions is conditioned on the preference, during the optimization process, we randomly sample the preferences in each iteration and optimize the acquisition function to generate the next candidate solutions for evaluation.
Preference-free
To marginalize the preferences, we leverage a key theoretical insight of the popular used hypervolume metric: when employing an infinite set of reference vectors uniformly sampled from the unit hypersphere $S=\{\boldsymbol{w}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{w}\|_{2}=1\}$, the hypervolume metric can be equivalently represented as the mean of a scalarization function (Deng and Zhang, 2019; Zhang and Golovin, 2020). Motivated by this, we firstly develop a preference-free acquisition function that utilizes the UCB utility defined in Equation (7) to approximate the uncertainty-aware hypervolume improvement (UHVI) for any candidate $\boldsymbol{x}$, given the evaluated inputs $\boldsymbol{X}=\{\boldsymbol{x}_{i}\}_{i=1}^{n}$:
$\alpha_{\text{UHVI}}(\boldsymbol{x})$ $=c_{m}\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[(c_{\boldsymbol{\lambda}})^{m}\max\{0,$ (8)
$[\min_{\boldsymbol{x}\in\boldsymbol{X}}\{-g_{\boldsymbol{\lambda}}(\boldsymbol{x})\}]^{m}-[-\alpha_{\text{UCB}}(\boldsymbol{x};\boldsymbol{\lambda})]^{m}\}],$
where $c_{m}=\frac{\pi^{m/2}}{2^{m}\Gamma(m/2+1)}$ is a constant that depends only on $m$. $c_{\boldsymbol{\lambda}}=\sqrt{\sum_{j=1}^{m}\frac{1}{\lambda_{j}^{2}}}$ is a transformation constant that depends only on $\boldsymbol{\lambda}$. The detailed derivations can be found in Appendix A.3.
Analogous to the derivation of UHVI, we draw upon the R2 indicator (Hansen and Jaszkiewicz, 1994; Brockhoff et al., 2012)—defined as the expected utility across a set of scalarization functions—to develop another preference-free acquisition function. This formulation, termed UR2I (uncertainty-aware R2 improvement), is designed to approximate the R2 improvement while explicitly accounting for predictive uncertainty:
$\alpha_{\text{UR2I}}(\boldsymbol{x})$ $=\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[\max\{0,$ (9)
$\min_{\boldsymbol{x}\in\boldsymbol{X}}\{-g_{\boldsymbol{\lambda}}(\boldsymbol{x})\}-[-\alpha_{\text{UCB}}(\boldsymbol{x};\boldsymbol{\lambda})]\}].$
5 Experiments
Baseline Algorithms
We compare our method against a wide range of classic and state-of-the-art MOBO algorithms. For consistency, MOEA/D-EGO (Zhang et al., 2009), TSEMO (Bradford et al., 2018), USEMO-EI (Belakaria et al., 2020), and DGEMO (Konakovic Lukovic et al., 2020) are all implemented using the DGEMO codebase based on Pymoo (Blank and Deb, 2020). In addition, we evaluate qNEHVI (Daulton et al., 2021), qParEGO (Daulton et al., 2020), and JES (Tu et al., 2022; Hvarfner et al., 2022) using the BoTorch framework (Balandat et al., 2020). Finally, we include the recently proposed BOFormer (Hung et al., 2025) as a leaning-based baseline algorithms, given its close relevance to our work.
Benchmark Problems
We first evaluate the performance of all algorithms on 6 synthetic functions featuring a wide spectrum of landscapes, following the experimental setup in BOFormer (Hung et al., 2025). Subsequently, we conduct extensive experiments on 10 real-world engineering design problems from the RE benchmark suite (Tanabe and Ishibuchi, 2020), which encompass diverse decision and objective space dimensions. Finally, we assess the methods on 4 hyperparameter optimization (HPO) tasks in a 3D object reconstruction problem (Mildenhall et al., 2021). Details of test problems are provided in Appendix B.1.
Experiment Setting
For all evaluated algorithms, each experiment is initialized with $2(d+1)$ randomly generated samples, followed by an optimization budget of 100 evaluations. To assess the quality of the Pareto front approximations produced by different algorithms, we employ the hypervolume (HV) indicator (Zitzler and Thiele, 2002) as the performance metric, given its strict monotonicity with respect to Pareto dominance. All experiments are conducted over 10 independent runs, we report the mean performance alongside the standard deviation with 95% confidence. Further technical details regarding the experimental configurations are provided in Appendix B.2 and B.3.
5.1 Optimization Performance
Table 1: Means (stds) of hypervolume metrics ($\uparrow$) on synthetic black-box functions over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm BC AR ABC ABD ACD BCD
MOEA/D-EGO 7.015e-01(7.5e-03) 1.203e+00(2.4e-03) 4.826e-01(8.1e-03) 6.934e-01(1.1e-02) 9.320e-01(1.4e-02) 5.735e-01(1.9e-02)
TSEMO 8.056e-01(9.1e-04) 1.172e+00(1.4e-02) 4.853e-01(4.5e-03) 7.114e-01(9.4e-03) 9.117e-01(2.6e-02) 6.938e-01(1.1e-02)
USEMO-EI 7.750e-01(1.2e-02) 1.118e+00(2.3e-02) 4.648e-01(8.2e-03) 7.298e-01(1.1e-02) 8.991e-01(2.5e-02) 6.766e-01(9.5e-03)
DGEMO 8.087e-01(1.3e-03) 1.158e+00(6.8e-02) 4.477e-01(3.8e-02) 5.918e-01(6.8e-02) 7.266e-01(9.1e-02) 7.054e-01(3.9e-03)
qNEHVI 7.962e-01(1.7e-03) 1.159e+00(1.8e-02) 4.929e-01(6.5e-03) 7.181e-01(1.0e-02) 9.167e-01(1.3e-02) 6.858e-01(1.5e-03)
qParEGO 7.843e-01(4.2e-03) 1.194e+00(4.4e-03) 4.782e-01(7.5e-03) 7.032e-01(9.1e-03) 9.285e-01(1.3e-02) 7.041e-01(4.5e-03)
JES 7.765e-01(3.1e-03) 1.086e+00(3.4e-02) 4.559e-01(1.1e-02) 6.598e-01(1.3e-02) 8.802e-01(2.8e-02) 6.947e-01(3.5e-03)
BOFormer 7.187e-01(1.4e-02) 1.092e+00(3.4e-02) 4.082e-01(1.7e-02) 6.383e-01(2.5e-02) 7.988e-01(3.8e-02) 6.715e-01(1.5e-02)
FoMEMO-EI 8.066e-01(6.2e-04) 1.192e+00(1.0e-02) 4.993e-01(1.2e-02) 7.308e-01(7.9e-03) 9.291e-01(1.5e-02) 7.346e-01(2.3e-03)
FoMEMO-UCB 8.075e-01(5.8e-04) 1.193e+00(8.4e-03) 4.878e-01(1.4e-02) 7.369e-01(8.5e-03) 9.407e-01(1.5e-02) 7.294e-01(6.2e-03)
FoMEMO-UHVI 8.052e-01(4.9e-04) 1.195e+00(6.3e-03) 4.908e-01(7.4e-03) 7.430e-01(8.4e-03) 9.519e-01(1.2e-02) 7.483e-01(7.6e-04)
FoMEMO-UR2I 8.016e-01(1.7e-03) 1.164e+00(2.2e-02) 4.973e-01(6.0e-03) 7.126e-01(9.1e-03) 9.256e-01(1.8e-02) 7.303e-01(1.2e-03)
Table 2: Means (stds) of hypervolume metrics ($\uparrow$) on engineering design problems over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm Four Bar Truss Pressure Vessel Hatch Cover Disc Brake Speed Reducer Gear Train Rocket Injector
MOEA/D-EGO 7.462e-01(1.4e-02) 1.129e+00(8.8e-03) 1.119e+00(2.3e-02) 1.280e+00(1.4e-02) 1.261e+00(1.3e-02) 5.623e-01(1.3e-01) 7.539e-01(1.1e-02)
TSEMO 8.506e-01(4.7e-04) 1.099e+00(1.6e-02) 1.022e+00(3.5e-02) 1.285e+00(3.0e-02) 1.297e+00(1.3e-03) 8.092e-01(2.3e-02) 8.618e-01(7.5e-04)
USEMO-EI 8.232e-01(4.3e-03) 1.153e+00(1.6e-03) 1.145e+00(6.2e-03) 1.305e+00(2.3e-03) 1.286e+00(3.7e-03) 6.433e-01(4.0e-02) 8.316e-01(3.7e-03)
DGEMO 8.526e-01(9.8e-04) 1.132e+00(6.3e-03) 1.003e+00(7.9e-02) 1.286e+00(2.9e-02) 1.304e+00(1.6e-04) 8.859e-01(3.2e-02) 8.605e-01(1.9e-03)
qNEHVI 8.244e-01(3.8e-03) 1.069e+00(1.6e-02) 1.119e+00(1.3e-02) 1.192e+00(3.1e-02) 1.263e+00(7.9e-03) 8.387e-01(2.6e-02) 8.182e-01(5.0e-03)
qParEGO 7.917e-01(1.7e-02) 1.099e+00(1.5e-02) 1.144e+00(6.7e-03) 1.199e+00(7.4e-02) 1.215e+00(1.4e-02) 7.813e-01(3.1e-02) 6.700e-01(2.8e-02)
JES 7.918e-01(1.0e-02) 8.795e-01(1.3e-01) 1.150e+00(3.5e-03) 1.249e+00(2.5e-02) 1.206e+00(1.0e-02) 6.096e-01(4.5e-02) 6.917e-01(2.4e-02)
BOFormer 7.308e-01(1.4e-02) 3.094e-01(1.7e-01) 1.119e+00(1.5e-02) 1.166e+00(2.3e-02) 1.191e+00(1.0e-02) 3.818e-01(5.7e-02) 6.208e-01(2.2e-02)
FoMEMO-EI 8.667e-01(2.0e-03) 1.151e+00(1.8e-03) 1.162e+00(1.5e-03) 1.309e+00(2.0e-03) 1.298e+00(7.3e-04) 9.107e-01(8.1e-03) 8.506e-01(4.6e-03)
FoMEMO-UCB 8.722e-01(1.2e-03) 1.150e+00(1.8e-03) 1.163e+00(1.9e-03) 1.312e+00(3.1e-04) 1.299e+00(6.2e-04) 9.270e-01(4.5e-03) 8.666e-01(2.4e-03)
FoMEMO-UHVI 8.609e-01(1.3e-03) 1.022e+00(5.4e-02) 1.127e+00(1.1e-02) 1.273e+00(1.6e-02) 1.263e+00(3.7e-03) 8.573e-01(9.6e-03) 8.245e-01(7.3e-03)
FoMEMO-UR2I 8.741e-01(5.0e-04) 1.086e+00(2.3e-02) 1.131e+00(4.8e-03) 1.299e+00(7.2e-03) 1.285e+00(3.6e-03) 8.961e-01(1.1e-02) 8.693e-01(1.5e-03)
Table 3: Means (stds) of hypervolume metrics ($\uparrow$) on hyperparameter optimization problems over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm Lego Materials Mic Ship
MOEA/D-EGO 1.184e+00(1.1e-03) 1.166e+00(2.2e-03) 1.257e+00(4.9e-03) 1.321e+00(1.5e-03)
TSEMO 1.174e+00(3.3e-03) 1.160e+00(5.1e-03) 1.251e+00(4.8e-03) 1.315e+00(2.1e-03)
USEMO-EI 1.179e+00(1.6e-03) 1.167e+00(2.8e-03) 1.255e+00(5.4e-03) 1.315e+00(3.6e-03)
DGEMO 1.171e+00(5.0e-03) 1.172e+00(3.2e-03) 1.273e+00(2.1e-02) 1.313e+00(3.6e-03)
qNEHVI 1.147e+00(9.8e-03) 1.144e+00(1.1e-02) 1.198e+00(1.9e-02) 1.293e+00(6.6e-03)
qParEGO 1.137e+00(1.2e-02) 1.150e+00(3.3e-03) 1.191e+00(2.0e-02) 1.291e+00(7.6e-03)
JES 1.167e+00(6.6e-03) 1.156e+00(3.2e-03) 1.212e+00(1.3e-02) 1.303e+00(4.5e-03)
BOFormer 1.176e+00(3.1e-03) 1.161e+00(3.6e-03) 1.244e+00(6.3e-03) 1.307e+00(3.3e-03)
FoMEMO-EI 1.177e+00(5.3e-03) 1.173e+00(2.0e-03) 1.261e+00(2.1e-02) 1.315e+00(2.9e-03)
FoMEMO-UCB 1.178e+00(7.3e-03) 1.175e+00(1.9e-03) 1.262e+00(1.8e-02) 1.316e+00(2.3e-03)
FoMEMO-UHVI 1.177e+00(6.2e-03) 1.176e+00(2.4e-03) 1.271e+00(1.6e-02) 1.318e+00(1.9e-03)
FoMEMO-UR2I 1.182e+00(2.6e-03) 1.179e+00(2.2e-03) 1.274e+00(1.2e-02) 1.320e+00(2.1e-03)
Tables 1, 2, and 3 summarize the sequential optimization results across all test problems. The experimental results demonstrate that our framework with four acquisition functions achieves superior performance in the majority of cases. This underscores that by pre-training the foundation model on an extensive and diverse corpus of synthetic data, the model showcases robust adaptability and generalization for problems that it has not encountered during training.
The core intuition driving this advancement lies in the direct and high-fidelity modeling of the unknown aggregated posteriors, combined with the high efficiency of fast in-context optimization. Consequently, it avoids relying on restrictive approximation assumptions of aggregated distributions (like in MOEA/D-EGO) or making excessive trade-offs for computational speed (like in qNEHVI), both of which can introduce modeling inaccuracies and ultimately compromise optimization performance. We further investigate the model’s scalability across varying objective and decision space dimensions, with detailed results and analysis provided in Appendix D.1.
5.2 Parallel Scalability
In addition to sequential optimization, we further investigate the scalability of our method in the context of parallel optimization. For preference-based acquisition functions, we generate a batch of $q$ candidates by conditioning $q$ random preferences sampled in each iteration. For preference-free variants, since approximating HV or R2 improvement via a finite set of random preferences introduces inherent stochastic deviation, we leverage this characteristic by independently repeating UHVI or UR2I calculations $q$ times to produce $q$ candidates for evaluation. We evaluate this parallel scheme under batch sizes of $q=5$ and $q=10$. Our results, detailed in Appendix D.2, demonstrate that even with these straightforward parallelization strategies, our preference-based methods consistently outperform MOBO algorithms that rely on well-designed batch selection heuristics. This performance gap highlights the advantage of using a foundation model pre-trained on a large variety of synthetic landscapes for optimization. Moreover, it demonstrates that modeling the preference-wise aggregated posterior is inherently well-suited for parallelization, providing a more robust and natural framework for multi-objective optimization without sophisticated batch selection processes.
5.3 Runtime Analysis
Table 4: Comparison of average query time per iteration for all algorithms in sequential optimization ($q=1$). The best, second-best, and third-best mean results for each case are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm 2 Objectives 3 Objectives 4 Objectives 6 Objectives
MOEA/D-EGO 38.6409 46.8428 49.9341 54.1983
TSEMO 5.6616 6.1670 5.8793 6.3524
USEMO-EI 8.1582 6.6809 7.2214 6.7517
DGEMO 36.2580 95.3770 N/A N/A
qParEGO 11.9262 19.0253 85.5916 185.4954
qNEHVI 10.1029 22.0746 62.9022 78.2812
JES 10.3307 33.4365 41.6235 44.9701
BOFormer 0.6679 0.7483 N/A N/A
FoMEMO-EI 2.5706 3.9841 3.1434 2.8501
FoMEMO-UCB 1.7201 1.1666 1.1155 2.0801
FoMEMO-UHVI 0.6189 0.6094 0.6197 0.7432
FoMEMO-UR2I 1.6248 1.7077 1.3686 1.1716
We evaluate the computational efficiency in sequential optimization across 4 engineering design problems (four bar truss, speed reducer, car side impact design and water resource planning problems) with objective counts ranging from 2 to 6, as shown in Table 4. The reported query time is averaged from 10 independent runs with 100 iterations each. Beyond the inherent time advantage of bypassing model retraining, here we ensure a rigorous comparison by monitoring only the duration required for acquisition function optimization to generate candidates, excluding the time allocated to model retraining (where applicable) and true function evaluations. This allows us to isolate and highlight the efficiency gains provided by the foundation model’s in-context optimization capabilities.
Due to the inference overhead of GP models, the computational cost of associated operations—such as the population-based optimization in MOEA/D-EGO, the intricate batch selection in DGEMO, and the expensive HV calculations in qNEHVI—increases drastically with problem complexity. In contrast, benefiting from the fast in-context learning of our foundation model, FoMEMO demonstrates superior time efficiency in optimizing acquisition functions as the objective dimension increases, some even achieving speedups of several orders of magnitude in high-dimensional scenarios. While BOFormer remains competitive in 2D and 3D objectives, it is still surpassed by FoMEMO-UHVI. Although the true function evaluations often dominate the model inference time in real-world problems, these results highlight the substantial efficiency gains provided by the foundation model paradigm, offering a more robust and scalable framework for many real-world challenges.
5.4 Modeling Accuracy Analysis
Since the efficacy of our method relies on modeling fidelity, in addition to providing a posterior visualization of a simple problem (see Appendix A.2), we further conduct a statistical analysis of the model’s ability to fit aggregation targets across problems with varying objective dimensions. Specifically, we monitor the accuracy of the predicted means throughout the training process by partitioning 1000 randomly generated datasets into training and testing sets using a 7:3 ratio. For each of the five independent experimental trials, a preference vector is randomly sampled to generate the aggregation targets. We employe RMSE and MAE metrics to quantify point-wise accuracy, alongside the $R^{2}$ score to evaluate the model’s capacity to capture the underlying data trends. We observe that as training progresses, the model’s predictive accuracy on unseen problems improves consistently. This evolution in modeling fidelity provides a fundamental explanation for the superior optimization performance demonstrated by our method.
Figure 2: The evolution of modeling accuracy in terms of regression error metrics (RMSE, MAE) and $R^{2}$ score during the training process.
6 Conclusion and Future Work
We propose a novel paradigm named FoMEMO, which, to the best of our knowledge, represents the first attempt to establish a foundation model to address a wide range of expensive MOPs in an in-context manner. Extensive experiments verify that our method exhibits remarkable adaptability and generalization in various synthetic problems and real-world applications, showing superior performance and high efficiency in most cases compared to existing methods.
In the future, we expect that this work opens up opportunities for many exciting directions worth exploring: (i) Develop architectures that allow for datasets with larger sizes, including embedding techniques and training schemes that efficiently handle higher-dimensional feature and objective spaces. (ii) Develop efficient acquisition functions tailored for high-dimensional and high-throughput parallel scenarios. (iii) Investigate whether fine-tuning can help improve the performance in domain-specific tasks.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Model and Algorithm Details
A.1 Details of Foundation Model and Pre-training
Table 5: Detailed information for the foundation model and pre-training process, including the hyperparameters and corresponding values of the model architecture, data generation and model training.
Type Name Value
Model Architecture Embedding Size 512
Hidden Size 1024
Attention Heads 4
Transformer Layers 12
Output Logits Size 1000
Data Generation GP Mean Prior 0.0
GP Kernel RBF
Output Scale 1.0
Output Noise $\mathcal{N}(0,10^{-4})$
Length Scale Prior $\Gamma(\alpha=3.0,\beta=6.0)$
Maximum Feature Dimension 30
Maximum Objective Dimension 6
Maximum Sample Length 256
Model Training Batch Size 128
Steps per Epoch 2048
Linear-warmup Epochs 5
Total Epochs 500
Learning Rate $5\times 10^{-5}$
Our foundational model mainly employs the recently proposed Prior-Data Fitted Networks (PFNs) (Müller et al., 2021) as the basic architecture, which is based on a Transformer encoder (Vaswani et al., 2017). We utilize linear layers to encode trajectory and query points, along with sampled preferences, as contextual inputs for the foundation model. Following the methods outlined in (Hollmann et al., 2022) and (Müller et al., 2023), we set a maximum dimension $K$ for encoding the inputs, which are applicable to both feature and objective dimensions. We zero-pad the sampled dimension $k$ and linearly scale the dimension by a factor of $K/k$ to ensure that the magnitudes of the input encodings are consistent across different dimensions. Since different query points should not depend on one another, an attention mask is implemented to allow each trajectory point can only attend to each other, while the query points can only attend to the trajectory points. To maintain the permutation invariance of the input dataset, the positional encodings are removed and the input points are fed to the model as the sum of their encoded embeddings. During training, we maintain a maximum sample length of $N$ with $n$ trajectory points and NnN-n query points, as used in (Müller et al., 2021). We sample each $n$ from $\{1,...,N-1\}$ with the weight of 1/(Nn)1/(N-n) to simulate varying trajectory sizes during the optimization process, while ensuring that each trajectory size has approximately the same number of query points for training. To model the continuous distribution for prediction, the regression head of PFNs discretizes the continuous distribution into a piecewise constant distribution, referred to as the Riemann distribution (Müller et al., 2021). This helps transform the model training from a regression task to a classification problem. The boundaries of each discrete interval are estimated by sampling a large synthetic dataset from the same synthetic prior, ensuring that each interval maintains equal probability in the prior data. Additionally, the last bar on each side is replaced with a suitably scaled half-normal distribution to enhance training stability. To account for the distributional shifts in aggregated posteriors across varying objective dimensions, we propose an objective-aware regression head. This architecture dynamically routes latent representations to specialized regressor conditioned on the specific objective dimension of the problem. By leveraging this dimension-specific mapping, the model can more accurately capture the differences of the aggregated posteriors with varying objective dimensions, thereby significantly bolstering predictive precision when generalizing to unseen optimization landscapes across diverse objective dimensions.
During the pre-training phase, we extensively sample from a data generation mechanism to create a large and diverse set of synthetic data. Initially, we sample the number of features $d$, the number of objectives $m$, and the trajectory length $n$ from 1 up to their predetermined maximum values. We then sample $m$ independent latent functions from a Gaussian Process (GP) prior with zero mean and the radial basis function (RBF) kernel:
$f\sim\mathcal{G}\mathcal{P}(0,k_{\textnormal{RBF}}(\cdot,\cdot)),\quad k_{\textnormal{RBF}}(\boldsymbol{x},\boldsymbol{x}^{\prime})=\sigma^{2}\exp(-\frac{\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|^{2}}{2l^{2}}),$ (10)
where $\sigma$ and $l$ represent the output scale and length scale kernel parameters, respectively. We also sample the length scale parameter for each feature dimension during the latent function sampling, along with an random observation noise to enhance function diversity. In the third step, we randomly sample input features from the $d$-dimensional unit cube, and then evaluate the sampled latent functions to generate the corresponding observations, resulting in $n$ trajectory samples and NnN-n query samples. After that, we randomly sample the artificial preferences from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$ to derive the aggregation targets $g$ given query inputs and preferences. The aggregation targets are then masked to formulate a supervised prediction problem, enabling the model to learn to predict the preference-wise aggregated posterior distribution $q_{\boldsymbol{\theta}}(g|\boldsymbol{x},D_{n};\boldsymbol{\lambda})$. We continuously synthesize data through extensive sampling and optimize the model parameters to minimize the cross-entropy loss function defined in Equation (5), which is equivalent to the expected KL divergence between the true posterior distribution and its approximation, up to an additive constant (Müller et al., 2021). The detailed information and hyperparameter settings regarding the foundational model and the pre-training process can be found in Table 5.
A.2 Visualization of Aggregated Posteriors
Figure 3: Details of the normalized objective functions for the ToyRobust problem, the aggregated posteriors (green solid lines with shaded regions) conditioned on 7 observations and 5 preferences, and the corresponding UCB acquisition functions (red solid lines).
To facilitate visualization, we employ a one-dimensional ToyRobust problem (Daulton et al., 2022) as a representative case to demonstrate the foundation model’s ability to predict aggregated posteriors with unseen objective functions, as illustrated in Figure 3. The results indicate that the model provides reliable posterior estimates directly from the evaluated trajectory of the unknown problem, without requiring further training or parameter updates. This underscores the model’s strong generalization capabilities beyond its training distribution. Specifically, the predicted mean accurately captures the global trends of individual sub-problems and also maintains similar local characteristics defined by adjacent preferences. Furthermore, the model yields well-calibrated uncertainty estimates, enabling acquisition functions—such as the Upper Confidence Bound (UCB) shown in the figure—to effectively balance exploration and exploitation throughout the optimization process.
A.3 Derivation of UHVI Acquisition Function
Definitions and preliminaries.
The derivation mainly depends on two types of definitions on Tchebycheff scalarizations and the theorem on hypervolume scalarization. We consider the following multi-objective optimization problem (MOP):
$\min_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{f}(\boldsymbol{x})=(f_{1}(\boldsymbol{x}),f_{2}(\boldsymbol{x}),\cdots,f_{m}(\boldsymbol{x})).$ (11)
We denote $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$ be a set of $n$ evaluated samples, where each $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{d})^{T}$ is the feature vector in the decision space $\mathcal{X}\subset\mathbb{R}^{d}$, and each $\boldsymbol{y}=(y_{1},y_{2},\ldots,y_{m})^{T}$ is the observation vector of the true objective function $\boldsymbol{f}(\boldsymbol{x})$. We denote $\boldsymbol{z}^{*}=(z_{1}^{*},\cdots,z_{m}^{*})^{T}$ as the ideal point (lower bound) in the objective space. We have the Tchebycheff scalarization function:
$s_{\boldsymbol{\lambda}}^{\text{tch}}(\boldsymbol{x})=\max_{1\leq j\leq m}\{\lambda_{j}(y_{j}(\boldsymbol{x})-z_{j}^{*})\},$ (12)
where $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\ldots,\lambda_{m})^{T}$ represents a preference sampled from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$, we can identify each Pareto optimal solution by minimizing $s_{\boldsymbol{\lambda}}^{\text{tch}}(\boldsymbol{x})$ with a specific but unknown preference (Choo and Atkins, 1983).
The Tchebycheff scalarization function can be transformed in another formulation with respect to a set of reference vectors $\boldsymbol{w}=(w_{1},w_{2},\ldots,w_{m})^{T}$ sampled from $S=\{\boldsymbol{w}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{w}\|_{2}=1\}$:
$s_{\boldsymbol{w}}^{\text{tch}}(\boldsymbol{x})=\max_{1\leq j\leq m}\{(y_{j}(\boldsymbol{x})-z_{j}^{*})/w_{j}\}.$ (13)
Leveraging Equation (13), we have the following theorem, which will help us approximate the hypervolume indicator and further derive UHVI acquisition function.
Theorem A.1 (Hypervolume Scalarization).
Let $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$ be a set of $n$ evaluated samples, $\boldsymbol{z}^{*}=(z_{1}^{*},\cdots,z_{m}^{*})^{T}$ is the ideal point, $\boldsymbol{r}=(r_{1},\cdots,r_{m})^{T}$ is a pre-defined reference point. The hypervolume of the evaluated samples can be expressed as an expectation of Tchebycheff scalarization function over a set of uniformly sampled reference vectors $S=\{\boldsymbol{w}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{w}\|_{2}=1\}$ (Deng and Zhang, 2019; Zhang and Golovin, 2020; Zhang et al., 2023):
$\textnormal{HV}=\prod_{j=1}^{m}(r_{j}-z_{j}^{*})-c_{m}\mathbb{E}_{\boldsymbol{w}\sim S}\{[\min_{1\leq i\leq n}s_{\boldsymbol{w}}^{\textnormal{tch}}(\boldsymbol{x}_{i})]^{m}\},$ (14)
where $c_{m}=\frac{\pi^{m/2}}{2^{m}\Gamma(m/2+1)}$ is a constant that depends only on $m$.
Utilizing the hypervolume scalarization and Equation (12), we can naturally derive the utility of UHVI acquisition function:
Lemma A.2 (Uncertainty-aware Hypervolume Improvement).
Given the distribution of the aggregation target $g_{\boldsymbol{\lambda}}(\cdot)=-s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\cdot)$ conditioned on evaluated samples $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$, the utility of hypervolume improvement in the normalized objective space at any query point $\boldsymbol{x}$ can be estimated by the following expectation:
$\alpha_{\textnormal{UHVI}}(\boldsymbol{x})=c_{m}\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[(c_{\boldsymbol{\lambda}})^{m}\max\{0,[\min_{1\leq i\leq n}\{-g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{i})\}]^{m}-[-\textnormal{UCB}(g_{\boldsymbol{\lambda}}(\boldsymbol{x}))]^{m}\}],$ (15)
where $c_{m}=\frac{\pi^{m/2}}{2^{m}\Gamma(m/2+1)}$ is a constant that depends only on $m$, $c_{\boldsymbol{\lambda}}=\sqrt{\sum_{j=1}^{m}\frac{1}{\lambda_{j}^{2}}}$ is a transformation constant that depends only on $\boldsymbol{\lambda}$, $\textnormal{UCB}(\cdot)$ denotes the upper confidence bound (UCB) metric.
Proof.
From Equation (14), given an updated evaluated samples $D_{n+1}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n+1}$ with newly added sample $\{\boldsymbol{x}_{n+1},\boldsymbol{y}_{n+1}\}$, the hypervolume can be represented as:
$\textnormal{HV}^{\prime}=\prod_{j=1}^{m}(r_{j}-z_{j}^{*})-c_{m}\mathbb{E}_{\boldsymbol{w}\sim S}\{[\min_{1\leq i\leq n+1}s_{\boldsymbol{w}}^{\textnormal{tch}}(\boldsymbol{x}_{i})]^{m}\},$ (16)
which induces the hypervolume improvement:
$\textnormal{HVI}(\boldsymbol{x}_{n+1})$ $=\textnormal{HV}^{\prime}-\textnormal{HV}$ (17)
$=c_{m}\mathbb{E}_{\boldsymbol{w}\sim S}\{[\min_{1\leq i\leq n}s_{\boldsymbol{w}}^{\textnormal{tch}}(\boldsymbol{x}_{i})]^{m}-[\min_{1\leq i\leq n+1}s_{\boldsymbol{w}}^{\textnormal{tch}}(\boldsymbol{x}_{i})]^{m}\}$
$=c_{m}\mathbb{E}_{\boldsymbol{w}\sim S}\{[\min_{1\leq i\leq n}s_{\boldsymbol{w}}^{\textnormal{tch}}(\boldsymbol{x}_{i})]^{m}-[\min\{\min_{1\leq i\leq n}s_{\boldsymbol{w}}^{\textnormal{tch}}(\boldsymbol{x}_{i}),s_{\boldsymbol{w}}^{\textnormal{tch}}(\boldsymbol{x}_{n+1})\}]^{m}\}$
$=c_{m}\mathbb{E}_{\boldsymbol{w}\sim S}[\max\{0,[\min_{1\leq i\leq n}s_{\boldsymbol{w}}^{\textnormal{tch}}(\boldsymbol{x}_{i})]^{m}-[s_{\boldsymbol{w}}^{\textnormal{tch}}(\boldsymbol{x}_{n+1})]^{m}\}].$
Given the transformation constant $c_{\boldsymbol{\lambda}}=\sqrt{\sum_{j=1}^{m}\frac{1}{\lambda_{j}^{2}}}$, each reference vector $\boldsymbol{w}$ can be uniquely represented by the preference $\boldsymbol{\lambda}$ by using $w_{j}=1/({\lambda_{j}c_{\boldsymbol{\lambda}})}$, the hypervolume improvement can be represented as:
$\textnormal{HVI}(\boldsymbol{x}_{n+1})=c_{m}\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[(c_{\boldsymbol{\lambda}})^{m}\max\{0,[\min_{1\leq i\leq n}s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\boldsymbol{x}_{i})]^{m}-[s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\boldsymbol{x}_{n+1})]^{m}\}].$ (18)
Our model predict the distribution of the aggregation target $g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{i})=-s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\boldsymbol{x}_{i})$ given each $\boldsymbol{x}_{i}$. To ensure that objectives of varying magnitudes share the same scale, we normalize the objectives to the range of [0,1] during the optimization process. Consequently, the hypervolume improvement in the normalized objective space can be reformulated as follows:
$\textnormal{HVI}(\boldsymbol{x}_{n+1})=c_{m}\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[(c_{\boldsymbol{\lambda}})^{m}\max\{0,[\min_{1\leq i\leq n}\{-g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{i})\}]^{m}-[-g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{n+1})]^{m}\}].$ (19)
Here the aggregation target $g_{\boldsymbol{\lambda}}(\cdot)$ is constrained to be non-positive after the objective normalization (see Figure 3 for an example). Since we wish to maximize the aggregation target at the query point $\boldsymbol{x}_{n+1}$ for maximizing hypervolume improvement, we replace the prediction $g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{n+1})$ with the $\textnormal{UCB}(\cdot)$ metric as an upper bound estimation, which introduces the calibrated uncertainty provided by our model for searching the candidates with a good balance of exploration and exploitation:
$\hat{\textnormal{HVI}}(\boldsymbol{x}_{n+1})=c_{m}\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[(c_{\boldsymbol{\lambda}})^{m}\max\{0,[\min_{1\leq i\leq n}\{-g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{i})\}]^{m}-[-\textnormal{UCB}(g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{n+1}))]^{m}\}].$ (20)
We can then derive the utility of UHVI acquisition function as $\boldsymbol{x}_{n+1}$ is arbitrary:
$\alpha_{\textnormal{UHVI}}(\boldsymbol{x})=c_{m}\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[(c_{\boldsymbol{\lambda}})^{m}\max\{0,[\min_{1\leq i\leq n}\{-g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{i})\}]^{m}-[-\textnormal{UCB}(g_{\boldsymbol{\lambda}}(\boldsymbol{x}))]^{m}\}],$ (21)
which leads to our claim in the lemma. ∎
Intuitively, maximizing $\alpha_{\textnormal{UHVI}}$ utility is equivalent to maximizing the predicted upper bounds of aggregation targets across all preference simultaneously. We find that our utility exhibits similar expression to that presented in (Song et al., 2024b). However, the key difference is that the formulation of (Song et al., 2024b) is based on the Gaussian process (GP) posterior for each objective, whereas our acquisition function is directly derived from the infinite preference-wise aggregated posteriors, which is intractable for traditional GP training and updates.
A.4 Derivation of UR2I Acquisition Function
Theorem A.3 (R2 Indicator).
Let $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$ be a set of $n$ evaluated samples, the R2 indicator of the evaluated samples is defined as an expectation of Tchebycheff scalarization function over a set of preferences sampled from the simplex $\Delta=\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}\mid\|\boldsymbol{\lambda}\|_{1}=1\}$ (Hansen and Jaszkiewicz, 1994; Brockhoff et al., 2012; Zhao et al., 2024):
$\textnormal{R}_{2}=\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[\min_{1\leq i\leq n}s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\boldsymbol{x}_{i})].$ (22)
Analogous to the derivation of the UHVI acquisition function presented above, the utility for the UR2I acquisition function can be naturally derived as follows. For the sake of brevity, the formal proof is omitted here.
Lemma A.4 (Uncertainty-aware R2 Improvement).
Given the distribution of the aggregation target $g_{\boldsymbol{\lambda}}(\cdot)=-s_{\boldsymbol{\lambda}}^{\textnormal{tch}}(\cdot)$ conditioned on evaluated samples $D_{n}=\{\boldsymbol{x}_{i},\boldsymbol{y}_{i}\}_{i=1}^{n}$, the utility of R2 improvement in the normalized objective space at any query point $\boldsymbol{x}$ can be estimated by the following expectation:
$\alpha_{\textnormal{UR2I}}(\boldsymbol{x})=\mathbb{E}_{\boldsymbol{\lambda}\sim\Delta}[\max\{0,\min_{1\leq i\leq n}\{-g_{\boldsymbol{\lambda}}(\boldsymbol{x}_{i})\}-[-\textnormal{UCB}(g_{\boldsymbol{\lambda}}(\boldsymbol{x}))]\}].$ (23)
Appendix B Experimental Setup Details
B.1 Benchmark Problems
Table 6: Detailed information of test benchmarks.
Type Name Feature Dimension ($d$) Objective Dimension ($m$)
Synthetic Function BC (Branin-Currin) 2 2
AR (Ackley-Rosenbrock) 2 2
ABC (Ackley-Branin-Currin) 2 3
ABD (Ackley-Branin-Dixon) 2 3
ACD (Ackley-Currin-Dixon) 2 3
BCD (Branin-Currin-Dixon) 2 3
Engineering Design Four Bar Truss Design 4 2
Pressure Vessel Design 4 2
Hatch Cover Design 2 2
Disc Brake Design 4 3
Speed Reducer Design 7 3
Gear Train Design 4 3
Rocket Injector Design 4 3
Car Side Impact Design 7 4
Conceptual Marine Design 6 4
Water Resource Planning 3 6
Hyperparameter Optimization Lego 5 2
Materials 5 2
Mic 5 3
ship 5 3
To evaluate the performance and generalization of our method, we employ a diverse suite of synthetic and real-world benchmarks, including 6 synthetic functions, 10 real-world engineering design problems and 4 hyperparameter optimization tasks. Specifically, we firstly utilize a combination of synthetic black-box functions characterized by a wide spectrum of landscapes, including unimodal versus multimodal structures, smooth versus non-smooth surfaces, and challenges involving narrow valleys and complex variable coupling, we utilize the implementation in BOFormer (Hung et al., 2025). Secondly, we incorporate the RE benchmark suite (Tanabe and Ishibuchi, 2020), which comprises real-world multi-objective engineering design problems from various fields such as pressure vessel, speed reducer, and rocket injector design, covering multi-objective ($m$ = 2, 3) to many-objective ($m$ = 4, 6) scenarios. Detailed description and characterization of the objective functions for these problems can be found in (Tanabe and Ishibuchi, 2020). Finally, we validate our approach on a hyperparameter optimization (HPO) problem for 3D Gaussian Splatting (3DGS). This problem aims to identify the optimal configurations of the the recent state-of-the-art method 3DGS (Kerbl et al., 2023) in 3D object reconstruction task (Mildenhall et al., 2021). The optimization objectives encompass rendering quality (measured by PSNR), model size, and the total number of Gaussian primitives. To comprehensively assess performance, we conduct evaluations across four diverse reconstruction scenes: Lego, Materials, Mic, and Ship from (Mildenhall et al., 2021). The detailed information of the benchmarks can be referred to Table 6.
B.2 Algorithm Hyperparameters
Since qNEHVI, qParEGO, JES, and our proposed method are all implemented based on the BoTorch framework (Balandat et al., 2020), we maintain consistent optimization settings for their respective acquisition functions. Specifically, we employ the multi-start L-BFGS-B optimizer with 20 restarts seeded from 1024 pseudo-random samples through BoTorch’s initialization heuristic, consistent with the protocol in (Ament et al., 2023). To strike a balance between optimization accuracy and computational efficiency, we utilize a stochastic approximation for the UHVI and UR2I utilities by randomly sampling 100 preference vectors during each optimization step. For all other baselines, we adhere to the default parameter configurations provided in their original code implementations unless otherwise specified.
B.3 Performance Metrics
To rigorously evaluate the quality of the Pareto fronts obtained by different algorithms, we employ the widely-used hypervolume (HV) indicator (Zitzler and Thiele, 2002) as the performance metric. For both synthetic black-box functions and real-world hyperparameter optimization tasks, we adopt the evaluation protocol from BOFormer (Hung et al., 2025), in which the true objective values are normalized using the pre-defined bounds. Consequently, we set the reference point to $\boldsymbol{r}=[1.1,...,1.1]^{T}\in\mathbb{R}^{m}$ for HV computation in the normalized objective space. For real-world engineering design problems, we utilize the approximate Pareto fronts provided by (Tanabe and Ishibuchi, 2020) as the ground truth, which are generated with a large number of evaluations and have also been widely used in other related studies. To ensure fairness, we firstly generate the ideal point $\boldsymbol{y}_{\text{ideal}}$ and nadir point $\boldsymbol{y}_{\text{nadir}}$ from the approximate Pareto front of the tested problem:
$\boldsymbol{y}_{\text{ideal}}=(\min y_{1}(\boldsymbol{x}_{1}),\ldots,\min y_{m}(\boldsymbol{x}_{m}))^{T},$ (24)
$\boldsymbol{y}_{\text{nadir}}=(\max y_{1}(\boldsymbol{x}_{1}),\ldots,\max y_{m}(\boldsymbol{x}_{m}))^{T},$
$\forall\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{m}\in\mathcal{P}_{s},$
where $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{m}$ are solutions in approximate Pareto set $\mathcal{P}_{s}$ from the ground truth. After that, we normalize the resulting Pareto fronts obtained from the algorithms using the bounds of ideal point and nadir point. Finally, we also set the reference point as $\boldsymbol{r}=[1.1,...,1.1]^{T}\in\mathbb{R}^{m}$ to calculate the hypervolume indicator of the normalized Pareto front obtained by each algorithm.
Appendix C Analysis of Model Training and Ablation Study
In this section, we provide a more detailed analysis of the model training process. To further demonstrate the superiority of our selected model configuration, we compare the optimization results across different configurations. Here, “Base” refers to the model setup used throughout this paper, which employs an RBF kernel, an objective-aware regression head, and Tchebycheff aggregation. For ablation analysis, we retrain the model by altering one key component at a time while keeping the rest of the configurations unchanged. Specifically, we experiment with Matern32 and Matern52 kernels, a uniform regression head that is agnostic to the objective dimension, and a smoothed version of the Tchebycheff aggregation with a smoothing parameter of 0.05 (Lin et al., 2024b).
C.1 Training Loss
Figure 4: The mean loss convergences with different model configurations.
We monitor the convergences of the mean loss during training presented by model with different configurations, represented in Figure 4. Despite minor fluctuations, each loss curve shows that it has essentially converged. We find that using the RBF kernel to generate data helps achieve a lower loss level, demonstrating the effectiveness of the RBF kernel in exploring a wider range of diverse data. While whether or not a objective-aware regression head is used has little impact on the reduction of loss. Notably, using smoothed aggregation leads to faster convergence and lower loss, which is related to the numerical stability brought by smoothing operation (Lin et al., 2024b), and thus facilitates gradient-based optimization of model parameters. However, there is a gap between the model pre-training phase and the subsequent downstream optimization process using the pre-trained models, and the performance of loss reduction cannot directly determine the optimization performance of the algorithm, which we will discuss further later.
C.2 Time Usage During Model Training
Table 7: Averaged time usage per step during model training (in seconds).
Model Configuration Data Time Step Time Forward Time
Base 0.0474 0.0323 0.0065
Matern32 Kernel 0.0491 0.0325 0.0066
Matern52 Kernel 0.0489 0.0326 0.0067
Objective-Agnostic Head 0.0518 0.0325 0.0100
Smoothed Aggregation 0.0588 0.0382 0.0067
In our method, the data synthesis and model training are performed simultaneously. To evaluate the advantages of using synthetic data for pre-training and in-context optimization with a single neural network forward pass, we record the detailed information on averaged time for each step utilized during training, as illustrated in Table 7. Among them, “data time” refers to the time spent synthesizing data for each step, “forward time” indicates the duration spent on model inference to generate logits for loss calculation, and “step time” denotes the time taken to compute the loss and perform gradient backpropagation to optimize the model parameters. Statistics results indicate that for all model configurations, the model optimization and data synthesis account for the majority of pre-training time. Although a large amount of data is synthesized during training using our sampling mechanism, the time spent on data synthesis is comparable to the time dedicated to optimizing model parameters. This efficiency can be largely attributed to the effective computation of the decomposition-based aggregation targets from extensive synthetic functions, which does not impose an unacceptable burden on pre-training as would be the case with expensive experimental data from real-world problems. Furthermore, it is noteworthy that a single forward propagation through the neural network is quite fast, this efficiency helps explain the rapid in-context inference observed once the pre-training is completed. In addition, we find that our chosen model configuration had advantages in all types of time usage.
C.3 Ablation Study of Optimization Performance
Table 8: Means (stds) of hypervolume metrics ($\uparrow$) on representative problems over 10 independent runs. The better mean result for each algorithm compared to the base configuration is marked in dark green.
Model Configuration Algorithm BC ABD Four Bar Truss Gear Train Materials Mic
Base FoMEMO-EI 8.066e-01(6.2e-04) 7.308e-01(7.9e-03) 8.667e-01(2.0e-03) 9.107e-01(8.1e-03) 1.173e+00(2.0e-03) 1.261e+00(2.1e-02)
FoMEMO-UCB 8.075e-01(5.8e-04) 7.369e-01(8.5e-03) 8.722e-01(1.2e-03) 9.270e-01(4.5e-03) 1.175e+00(1.9e-03) 1.262e+00(1.8e-02)
FoMEMO-UHVI 8.052e-01(4.9e-04) 7.430e-01(8.4e-03) 8.609e-01(1.3e-03) 8.573e-01(9.6e-03) 1.176e+00(2.4e-03) 1.271e+00(1.6e-02)
FoMEMO-UR2I 8.016e-01(1.7e-03) 7.126e-01(9.1e-03) 8.741e-01(5.0e-04) 8.961e-01(1.1e-02) 1.179e+00(2.2e-03) 1.274e+00(1.2e-02)
Matern32 Kernel FoMEMO-EI 8.071e-01(5.0e-04) 7.354e-01(8.9e-03) 8.676e-01(1.1e-03) 9.134e-01(5.0e-03) 1.175e+00(2.0e-03) 1.263e+00(3.9e-03)
FoMEMO-UCB 8.063e-01(1.4e-03) 7.501e-01(1.1e-02) 8.693e-01(2.0e-03) 9.145e-01(8.4e-03) 1.175e+00(2.4e-03) 1.235e+00(2.6e-02)
FoMEMO-UHVI 7.841e-01(3.6e-02) 7.340e-01(7.8e-03) 8.560e-01(2.1e-03) 8.456e-01(9.8e-03) 1.175e+00(2.9e-03) 1.261e+00(2.1e-02)
FoMEMO-UR2I 8.018e-01(1.0e-03) 7.192e-01(8.6e-03) 8.697e-01(1.5e-03) 8.890e-01(7.8e-03) 1.174e+00(9.1e-03) 1.244e+00(1.6e-02)
Matern52 Kernel FoMEMO-EI 8.074e-01(6.0e-04) 7.430e-01(9.1e-03) 8.687e-01(9.9e-04) 9.090e-01(7.0e-03) 1.175e+00(2.6e-03) 1.254e+00(9.5e-03)
FoMEMO-UCB 8.083e-01(5.3e-04) 7.495e-01(9.1e-03) 8.736e-01(8.1e-04) 9.156e-01(6.5e-03) 1.177e+00(2.0e-03) 1.253e+00(2.0e-02)
FoMEMO-UHVI 8.074e-01(4.3e-04) 7.396e-01(7.6e-03) 8.610e-01(1.4e-03) 8.383e-01(8.0e-03) 1.175e+00(1.6e-03) 1.267e+00(1.4e-02)
FoMEMO-UR2I 7.975e-01(1.2e-03) 7.178e-01(1.1e-02) 8.753e-01(1.2e-03) 8.658e-01(1.1e-02) 1.178e+00(1.3e-03) 1.255e+00(6.7e-03)
Objective-Agnostic Head FoMEMO-EI 8.073e-01(9.1e-04) 7.192e-01(8.0e-03) 8.666e-01(1.9e-03) 9.105e-01(7.7e-03) 1.173e+00(2.5e-03) 1.261e+00(1.2e-02)
FoMEMO-UCB 8.085e-01(6.1e-04) 7.305e-01(9.2e-03) 8.721e-01(1.6e-03) 9.221e-01(4.6e-03) 1.177e+00(2.3e-03) 1.263e+00(1.2e-02)
FoMEMO-UHVI 8.070e-01(3.6e-04) 7.349e-01(1.1e-02) 8.639e-01(7.0e-04) 8.506e-01(9.9e-03) 1.174e+00(2.1e-03) 1.267e+00(1.6e-02)
FoMEMO-UR2I 8.069e-01(1.0e-03) 7.158e-01(6.8e-03) 8.769e-01(5.6e-04) 8.838e-01(6.1e-03) 1.176e+00(2.1e-03) 1.262e+00(9.1e-03)
Smoothed Aggregation FoMEMO-EI 7.995e-01(2.4e-03) 7.407e-01(7.0e-03) 8.629e-01(1.8e-03) 9.064e-01(7.5e-03) 1.176e+00(2.0e-03) 1.271e+00(1.6e-02)
FoMEMO-UCB 8.027e-01(1.4e-03) 7.397e-01(8.9e-03) 8.683e-01(1.8e-03) 9.279e-01(3.5e-03) 1.177e+00(1.7e-03) 1.269e+00(1.3e-02)
FoMEMO-UHVI 7.977e-01(1.4e-03) 7.280e-01(8.8e-03) 8.577e-01(3.3e-03) 8.329e-01(1.7e-02) 1.174e+00(1.9e-03) 1.266e+00(1.6e-02)
FoMEMO-UR2I 7.903e-01(1.9e-03) 6.897e-01(5.0e-02) 8.669e-01(7.7e-04) 8.857e-01(4.6e-03) 1.176e+00(1.9e-03) 1.267e+00(1.3e-02)
We further compared the optimization performance of all proposed acquisition functions with different model configurations on 6 representative problems, as shown in Table 8. For each algorithm with different model configurations, if it outperforms the corresponding algorithm in base configuration, we mark it in dark green to visualize the impact of different model configurations on the final optimization performance. The results show that our chosen model configuration performs better in most cases. We find that using Matern32 and Matern52 yield better results especially for synthetic problems involving Branin functions. The reason may lie that the RBF kernel, being infinitely differentiable, imposes a strong smoothness prior that hinders its ability to capture the localized transitions and high-curvature regions between the peaks and valleys of the Branin landscape. In contrast, the Matern class of kernels relaxes these smoothness constraints, which effectively balances the global trend capturing and local flexibility required to navigate the distinctive landscape of the Branin function. In addition, we observe that using the objective-agnostic head yields inferior results compared to the base configuration in most cases. This demonstrates that while loss gaps may be marginal during training, the adaptive selection of dimension-specific regressors during inference is pivotal for enhancing final optimization outcomes. Finally, empirical results indicate that the configuration with smoothed aggregation generally underperforms the baseline. This disparity likely stems from the fact that while smoothing facilitates model training, the approximation errors introduced by smoothing operation may bias the accurate fitting of the optimal region in the exact aggregation function (Lin et al., 2024b), further misleading the precise identification of the global optimum in the sub-problem during the optimization process.
Appendix D Analysis of Method Scalability
D.1 Dimensional Scalability
Table 9: Means (stds) of hypervolume metrics ($\uparrow$) on 3 engineering design problems with many objectives and a multi-modal WFG problem with complex search space over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm Car Side Impact (7d4m) Conceptual Marine (6d4m) Water Resource (3d6m) WFG (30d3m)
MOEA/D-EGO 7.039e-01(5.7e-03) 5.915e-01(2.4e-02) 1.412e+00(1.3e-02) 2.754e-01(5.3e-03)
TSEMO 6.942e-01(7.9e-03) 7.358e-01(7.9e-03) 1.198e+00(5.7e-02) 2.869e-01(4.3e-03)
USEMO-EI 7.151e-01(7.2e-03) 6.750e-01(1.1e-02) 1.433e+00(9.4e-03) 2.842e-01(5.3e-03)
DGEMO N/A N/A N/A 2.792e-01(4.6e-03)
qNEHVI 5.900e-01(7.0e-03) 6.472e-01(2.2e-02) 1.342e+00(4.8e-02) 3.276e-01(2.2e-02)
qParEGO 5.740e-01(2.3e-02) 3.887e-01(2.6e-02) 1.404e+00(3.7e-02) 3.213e-01(1.5e-02)
JES 5.629e-01(1.7e-02) 4.628e-01(2.9e-02) 1.296e+00(2.6e-02) 3.314e-01(7.5e-03)
BOFormer N/A N/A N/A 3.086e-01(5.2e-03)
FoMEMO-EI 7.538e-01(7.4e-03) 7.505e-01(1.9e-02) 1.470e+00(4.7e-03) 3.159e-01(4.3e-03)
FoMEMO-UCB 7.739e-01(8.1e-03) 7.814e-01(1.5e-02) 1.475e+00(3.9e-03) 3.150e-01(4.4e-03)
FoMEMO-UHVI 7.156e-01(4.7e-03) 6.012e-01(1.6e-02) 1.430e+00(1.9e-02) 3.173e-01(7.8e-03)
FoMEMO-UR2I 7.552e-01(5.1e-03) 7.587e-01(3.8e-03) 1.370e+00(2.0e-02) 3.397e-01(1.4e-02)
In this subsection, we further evaluate the dimensional scalability of our method, including both objective and feature dimensions. Firstly, we consider 3 more challenging real-world engineering design problems with higher objective dimensions, including car side impact design and conceptual marine design problems, both with 4 objectives, as well as water resource planning problem, which has 6 objectives (Tanabe and Ishibuchi, 2020). The experimental results are shown in Table 9. We find that our method demonstrates superior scalability and maintains robust performance in many-objective optimization scenarios. A primary factor contributing to this efficacy of our approach is its ability to directly predict the exact aggregated posterior as a universal basis, which avoids the performance-degrading compromises often found in existing methods. In contrast, traditional methods suffer from significant limitations as the objective dimensionality increases: qNEHVI faces the computational burden of non-dominated partitioning and hypervolume calculations, which become computationally intractable in high-dimensional spaces. MOEA/D-EGO introduces approximation errors due to the non-Gaussian nature of the max operation when using Gaussian distribution to estimate the unknown aggregation distribution in high-dimensional objective spaces. By bypassing these approximations, our method preserves modeling fidelity without sacrificing efficiency. Secondly, unlike qParEGO that model each sub-problem repeatedly, our method utilizes the advantage of similar patterns shared by sub-problems with adjacent preferences, which means the belief updates of each sub-problem during training or optimization can be efficiently facilitated by its neighbors.
To examine the scalability of our approach across varying feature dimensions, we employ the WFG4 benchmark (Huband et al., 2006), which allows for flexible expansion of the decision space. Characterized by its pronounced multi-modality, WFG4 presents a complex landscape that rigorously evaluates an algorithm’s ability to maintain global exploration and circumvent premature convergence. For this evaluation, we consider a 30-dimensional instance as a representative case. This setup aligns with the maximum dimensionality encountered during our current training phase, serving as a critical testbed to verify the model’s generalization robustness at the boundary of its training scope. Experimental results show that FoMEMO-UR2I consistently maintains superior performance even in this challenging scenario. While our architecture is readily adaptable to higher-dimensional spaces through expanded training, we focus on the aforementioned configurations here to establish a performance baseline for dimensionality scalability. Expanding the current framework to scale toward even more massive objective and decision spaces has been designated as a primary objective for our future work.
D.2 Parallel Scalability
In this subsection, we provide supplementary experimental results to further validate the parallel scalability of our approach. Tables 10 and 11 report the performance across all algorithms with batch sizes of $q=5$ and $q=10$ per iteration, respectively. To ensure a rigorous and consistent comparison, a total evaluation budget of 100 samples is allocated for each configuration, excluding the initial design points. Our findings reveal that even with a straightforward batch generation strategy, our method consistently outperforms methods such as TSEMO and DGEMO in most cases, despite their reliance on sophisticated batch selection heuristics. Notably, among our proposed variants, the two preference-based acquisition functions yield the most superior results. This performance gap underscores the advantage of leveraging a foundation model pre-trained on diverse synthetic landscapes. Furthermore, it demonstrates that modeling aggregated posteriors parameterized by preferences is inherently amenable to parallelization, providing a more robust and natural framework for multi-objective optimization without focusing on developing more sophisticated batch-selection strategies.
Table 10: Means (stds) of hypervolume metrics ($\uparrow$) in parallel optimization (q=5) on representative problems over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm BC ABD Four Bar Truss Gear Train Materials Mic
MOEA/D-EGO 7.052e-01(1.5e-02) 6.227e-01(5.3e-02) 7.658e-01(1.7e-02) 5.365e-01(1.2e-01) 1.164e+00(6.2e-03) 1.247e+00(9.0e-03)
TSEMO 8.081e-01(5.8e-04) 7.117e-01(1.5e-02) 8.520e-01(1.9e-03) 7.811e-01(3.8e-02) 1.155e+00(5.9e-03) 1.238e+00(7.5e-03)
USEMO-EI 7.882e-01(5.6e-03) 7.057e-01(1.8e-02) 8.287e-01(6.3e-03) 6.742e-01(4.9e-02) 1.150e+00(5.3e-03) 1.243e+00(1.5e-02)
DGEMO 8.068e-01(8.8e-04) 5.838e-01(5.9e-02) 8.549e-01(3.0e-03) 8.413e-01(6.8e-02) 1.162e+00(7.8e-03) 1.272e+00(1.8e-02)
qNEHVI 7.990e-01(1.5e-03) 7.105e-01(1.1e-02) 8.240e-01(5.8e-03) 7.810e-01(7.5e-02) 1.150e+00(7.9e-03) 1.232e+00(2.1e-02)
qParEGO 7.744e-01(3.3e-03) 6.852e-01(8.5e-03) 7.651e-01(1.8e-02) 6.832e-01(7.2e-02) 1.142e+00(1.6e-02) 1.176e+00(2.1e-02)
JES 7.575e-01(5.5e-03) 6.601e-01(1.6e-02) 7.861e-01(9.1e-03) 5.783e-01(7.0e-02) 1.157e+00(4.3e-03) 1.211e+00(2.1e-02)
FoMEMO-EI 8.064e-01(1.2e-03) 7.351e-01(8.0e-03) 8.656e-01(2.2e-03) 9.043e-01(7.1e-03) 1.171e+00(2.2e-03) 1.245e+00(1.4e-02)
FoMEMO-UCB 8.072e-01(5.3e-04) 7.311e-01(5.1e-03) 8.698e-01(1.5e-03) 9.229e-01(5.3e-03) 1.176e+00(2.4e-03) 1.236e+00(1.9e-02)
FoMEMO-UHVI 8.036e-01(6.4e-04) 7.298e-01(6.5e-03) 8.567e-01(1.7e-03) 8.381e-01(1.2e-02) 1.173e+00(1.5e-03) 1.250e+00(2.5e-02)
FoMEMO-UR2I 7.922e-01(1.8e-03) 6.538e-01(9.8e-03) 8.420e-01(4.5e-03) 8.288e-01(2.1e-02) 1.156e+00(4.5e-03) 1.214e+00(1.2e-02)
Table 11: Means (stds) of hypervolume metrics ($\uparrow$) in parallel optimization (q=10) on representative problems over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm BC ABD Four Bar Truss Gear Train Materials Mic
MOEA/D-EGO 7.221e-01(2.1e-02) 5.363e-01(6.9e-02) 7.872e-01(2.0e-02) 3.939e-01(1.3e-01) 1.154e+00(5.7e-03) 1.238e+00(1.1e-02)
TSEMO 8.074e-01(7.6e-04) 6.512e-01(3.3e-02) 8.521e-01(2.1e-03) 7.060e-01(6.9e-02) 1.145e+00(9.6e-03) 1.242e+00(9.8e-03)
USEMO-EI 7.708e-01(1.4e-02) 6.633e-01(2.2e-02) 8.267e-01(1.2e-02) 6.155e-01(9.7e-02) 1.143e+00(7.9e-03) 1.231e+00(1.4e-02)
DGEMO 8.063e-01(1.3e-03) 5.228e-01(5.8e-02) 8.713e-01(3.0e-03) 8.221e-01(5.7e-02) 1.149e+00(7.3e-03) 1.242e+00(1.3e-02)
qNEHVI 7.975e-01(2.4e-03) 7.222e-01(7.5e-03) 8.194e-01(3.9e-03) 7.600e-01(8.7e-02) 1.145e+00(9.2e-03) 1.207e+00(2.1e-02)
qParEGO 7.722e-01(7.8e-03) 6.892e-01(1.2e-02) 7.286e-01(2.3e-02) 6.207e-01(1.0e-01) 1.150e+00(1.1e-02) 1.175e+00(1.4e-02)
JES 7.622e-01(3.2e-03) 6.245e-01(2.1e-02) 7.813e-01(8.3e-03) 5.677e-01(6.8e-02) 1.159e+00(3.0e-03) 1.222e+00(1.6e-02)
FoMEMO-EI 8.063e-01(8.5e-04) 7.248e-01(7.4e-03) 8.652e-01(1.8e-03) 8.987e-01(1.5e-02) 1.164e+00(4.7e-03) 1.237e+00(2.7e-02)
FoMEMO-UCB 8.064e-01(6.5e-04) 7.347e-01(1.2e-02) 8.679e-01(2.8e-03) 9.188e-01(6.1e-03) 1.167e+00(5.4e-03) 1.247e+00(2.6e-02)
FoMEMO-UHVI 7.881e-01(2.1e-02) 7.179e-01(1.0e-02) 8.494e-01(1.7e-03) 8.385e-01(1.6e-02) 1.166e+00(4.8e-03) 1.259e+00(1.5e-02)
FoMEMO-UR2I 7.297e-01(2.5e-02) 5.862e-01(1.9e-02) 7.577e-01(1.7e-02) 6.736e-01(6.4e-02) 1.137e+00(1.2e-02) 1.180e+00(2.0e-02)
Appendix E Limitations and Potential Future Work
E.1 GP Priors Limitations
In this paper, we focus on addressing a general class of unconstrained multi-objective expensive optimization problems characterized by continuous, nonlinear objective functions that exhibit a degree of smoothness and allow for uncertainty quantification, which are prevalent in numerous real-world applications, making them well-suited for Gaussian Process (GP) modeling. Therefore, we can extensively sample from GP priors to generate a substantial amount of synthetic training data, simulating a wide range of potential mappings that may exist in such real-world scenarios. The experimental results further demonstrate the applicability and generalization of our method. However, when confronted with problems that involve special knowledge completely beyond GP priors, our method may struggle to generalize to these specific types of cases, as we cannot train the model indefinitely to traverse the infinite space of function hyperparameters. Potential solutions include incorporating expert knowledge to design specialized priors during synthetic data generation, or interacting with the evaluation environment during optimization to discover and enrich potential priors. Since synthetic data can be generated from the designed priors through extensive sampling, these enhancements can be seamlessly integrated into our existing framework. Furthermore, instead of permanently freezing the model parameters after a single round of synthetic pre-training, it may be advantageous to fine-tune the pre-trained model using a small amount of domain-specific data.
E.2 Scalability Challenges
Although we have evaluated the scalability of our approach across problems with varying feature and objective spaces (see Section D.1), those involving significantly higher-dimensional features or objectives pose considerable challenges. While our framework can easily accommodate higher-dimensional data during training, the sample complexity required to adequately represent such spaces scales exponentially, which imposes a prohibitive challenge to the synthesis of a representative pre-training corpus. This curse of dimensionality mirrors a fundamental bottleneck in the multi-objective optimization community, where the exponential expansion of the search space remains a primary challenge for both surrogate modeling and search efficiency. To this end, several promising avenues worth further exploration. In the context of high-dimensional feature spaces, the development of efficient dimensionality reduction methods and embedding techniques, along with effective search strategies such as the trust region Bayesian optimization (TuRBO) algorithm (Eriksson et al., 2019), presents significant potential. For problems involving a large number of objectives, the number of solutions required to approximate the entire Pareto set increases exponentially, resulting in significant overhead for both optimization and decision-making. Therefore, rather than pursuing a comprehensive set of Pareto solutions, recent work has focused on identifying a small set of representative solutions, ensuring that each objective can be well addressed by at least one solution within the set (Lin et al., 2024a; Liu et al., 2024). This motivates the development of foundational models and parallel acquisition functions based on the set scalarizations (Lin et al., 2024a).